# quantum neural networks, Bernstein’s inequality, neural tangent kernel, Clifford group

### v1

@@ Abstract
We propose an efficient classical algorithm to estimate the Neural Tangent Kernel (NTK) associated with a broad class of quantum neural networks. These networks consist of arbitrary unitary operators belonging to the Clifford group interleaved with parametric gates given by the time evolution generated by an arbitrary Hamiltonian belonging to the Pauli group. The proposed algorithm leverages a key insight: the average over the distribution of initialization parameters in the NTK definition can be exactly replaced by an average over just four discrete values, chosen such that the corresponding parametric gates are Clifford operations. This reduction enables an efficient classical simulation of the circuit. Combined with recent results establishing the equivalence between wide quantum neural networks and Gaussian processes [Girardi et al., Comm. Math. Phys. 406, 92 (2025); Melchor Hernandez et al., arXiv:2412.03182], our method enables efficient computation of the expected output of wide, trained quantum neural networks, and therefore shows that such networks cannot achieve quantum advantage.
@@ 1 Introduction
Quantum Machine Learning (QML) is an interdisciplinary field that merges the principles of quantum computing with classical machine learning techniques de2019primer ; schuld2015 ; pastorello2023concise . One of the core ideas of QML is to harness quantum algorithms and the unique properties of quantum mechanics such as superposition, entanglement, and quantum parallelism to enhance the performance of deep neural models biamonte2017 . Quantum neural networks constitute the quantum version of deep neural models. These new models are based on quantum circuits and generate functions given by the expectation values of a quantum observable measured on the output of a quantum circuit made by parametric gates girardi2024 ; schuld2018 . These parameters are typically optimized by gradient descent, which involves iterative adjustment to minimize a cost function and improve the performance of the quantum circuit in the processing and analysis of data schuld2021effect . Several works have focused on the analysis of quantum neural networks, as it is believed that they can combine the computational power of quantum computers with the capabilities of deep learning algorithms lloyd2020quantum . Ref. liu2021 shows that an exponential quantum speed-up can be obtained via the use of a quantum-enhanced feature space, where each data point is mapped in a non-linear way to a quantum state, and then classified by a linear classifier in a high-dimensional Hilbert space havlivcek2019 . Nevertheless, a significant disadvantage lies in the need to determine the appropriate parameters to configure the quantum circuit beforehand, and it is not yet clear whether these parameters can be effectively obtained using a variational technique cinelli2021var .
@@
A rigorous mathematical characterization of the training dynamics of quantum neural networks becomes possible in the limit of infinite width. Ref. abedi2023 proves that in the limit of infinite width with the depth kept fixed, trained quantum neural networks with constant depth operate in the lazy regime (i.e., the maximum amount by which the training can change a parameter tends to zero) and are capable of perfectly fitting the training examples. However, such a regime is not interesting for quantum advantage, since at finite depth, the past light-cone of each measured qubit has a finite size and the expectation value of the measured observable can be efficiently estimated on a classical computer. The recent works girardi2024 ; hernandez2024 consider the joint limit of infinite width and depth. Under the hypothesis that the depth grows at most logarithmically with respect to the number of qubits, they prove that the probability distribution of the trained model function converges in distribution to a Gaussian process. The key element of the proof is showing that the training happens in the lazy regime, and therefore the dependence of the model function on the parameters can be approximated by its linearized version near the initialization values. Consequently, in the limit the model becomes linear and the training has an analytic solution whose probability distribution is Gaussian with analytically computable mean and covariance. Such a linear model is governed by a kernel called the Neural Tangent Kernel (NTK), which has been investigated in several recent studies schuld2015 ; Shirai:2021jpy ; PRXQuantum.3.030323 ; Yu_2024 ; scala2025towards .
@@
Let us recall that, given a general parametric model $f_{\theta}:\mathcal{X}\rightarrow\mathbb{R}$, with $\theta\in\mathbb{R}^{L}$, where $\mathcal{X}$ is the feature space, the empirical NTK is defined as NEURIPS2018_5a4be1fa :
@@ 1.1 Our contribution
The key idea of the algorithm is that, since $f_{\theta}(x)$ is a quadratic polynomial in the entries of the unitary matrix associated to each gate and the analytic NTK is quadratic in $f$, the expectation value with respect to the uniform distribution of the parameters can be replaced by the expectation value with respect to the uniform distribution on the set $\left\{0,\frac{\pi}{2},\pi,\frac{3\pi}{2}\right\}$. For such values of the parameters, the parametric gates belong to the Clifford group, and since also the nonparametric gates belong to the Clifford group by hypothesis, the whole $U_{x,\theta}$ belongs to the Clifford group. Therefore, the analytic NTK can be computed via the efficient algorithm for Clifford quantum circuits Aaronson_2004 .
Let us now state informally our second contribution. The average of the Gaussian process associated to a trained quantum neural network has a limit $\mu_{\infty}:\mathcal{X}\to\mathbb{R}$ for infinite training time. Such a limit depends only on the NTK of the network and on the training data (see subsection 2.4 and girardi2024 ; hernandez2024 for the details). We can now leverage the algorithm of Theorem 1 to estimate $\mu_{\infty}$:
@@
    The output $\hat{\mu}_{\infty}$ of the algorithm above satisfies
    $\left|\hat{\mu}_{\infty}(x)-\mu_{\infty}(x)\right|<\epsilon\qquad\forall\,x\in\mathcal{X}$ (1.10)
@@
Our work conceptually aligns with a recent series of results angrisani2024 ; ReardonSmith2024improvedsimulation ; PRXQuantum.6.010337 ; martinez25 ; cerezo2024doesprovableabsencebarren demonstrating that a wide class of variational quantum circuits can be simulated efficiently by classical computers. In particular, Ref. angrisani2024 shows that, if each layer of the circuit contains random iid one-qubit gates sampled from the Haar distribution applied to all the qubits, then, regardless of how the two-qubit gates are placed, the untrained circuit can be simulated efficiently with high probability. On the contrary, our results require the validity of the linear approximation governed by the NTK, but are able to estimate the model function of the trained circuit.
@@ 2.2 The Pauli group and the Clifford group
For details about properties of these groups, see for instance mastel2023clifford .
@@ 2.4 The neural tangent kernel
Given any function $g:\mathcal{X}\to\mathbb{R}$, we will often use the following notation:
@@
We suppose that the expected value in (2.11) is computed with the entries of $\theta$ sampled from iid uniform distributions on $0,2\pi)$.
Assumption 2.
@@
The analytic NTK is the cornerstone in the analysis of the gradient flow equation
where $\mathcal{L}$} is the quadratic cost
and $\eta>0$ is the learning rate. The loss function $\mathcal{L}(\theta)$ is not convex girardi2024 , hence gradient flow is a priori not guaranteed to converge to a global minimum abedi2023 . From the gradient-flow equation (2.13) and the chain rule, one gets that
@@
where $\nabla_{\theta}f_{\theta_{t}}(X_{{\rm train}}^{T})$ denotes the gradient of $f_{\theta_{t}}(X_{{\rm train}}^{T})$ with respect to $\theta$ while $\nabla_{f_{\theta_{t}}(X_{{\rm train}})}{\mathcal{L}}(\theta_{t})$ indicates the gradient of the cost function $\mathcal{L}$} with respect to $f_{\theta_{t}}(X_{{\rm train}})$. Recall that T is the transposition operator. Let us also notice that (2.15) can be written as
$\displaystyle\begin{cases}\frac{{\rm d}\theta_{t}}{{\rm d}t}=-\eta\nabla_{\theta}f_{\theta_{t}}(X_{{\rm train}}^{T})\nabla_{f_{\theta_{t}}(X_{{\rm train}})}{\mathcal{L}}(\theta_{t}),\\ \frac{{\rm d}}{{\rm d}t}f_{\theta_{t}}(x)=-\eta\hat{K}_{\theta_{t}}(x,X_{{\rm train}}^{T})\nabla_{f_{\theta_{t}}(X_{{\rm train}})}{\mathcal{L}}(\theta_{t}).\end{cases}$ (2.16)
Refs. abedi2023 ; girardi2024 ; hernandez2024 consider the linear approximation of the function $f_{\theta_{t}}(\cdot)$ around the initialization values $\theta_{0}$ of the parameters:
$f_{\theta_{t}^{{\rm lin}}}^{{\rm lin}}(x)=f_{\theta_{0}}(x)+\nabla_{\Theta}f_{\theta_{0}}(x)^{T}(\theta_{t}^{{\rm lin}}-\theta_{0}).$ (2.17)
In this approximation, (2.16) becomes
$\displaystyle\begin{cases}&\frac{{\rm d}\theta_{t}^{{\rm lin}}}{{\rm d}t}=-\eta\nabla_{\theta}f_{\theta_{0}}(X_{{\rm train}}^{T})\nabla_{f_{\theta_{t}^{\rm lin}}^{{\rm lin}}(X_{{\rm train}})}{\mathcal{L}}^{{\rm lin}}(\theta_{t}^{{\rm lin}}),\\ &\frac{{\rm d}}{{\rm d}t}f_{\theta_{t}^{{\rm lin}}}^{{\rm lin}}(x)=-\eta\hat{K}_{\theta_{0}}(x,X_{{\rm train}}^{T})\nabla_{f_{\theta_{t}^{{\rm lin}}}^{{\rm lin}}(X_{{\rm train}})}{\mathcal{L}}^{{\rm lin}}(\theta_{t}^{{\rm lin}}),\end{cases}$ (2.18)
where ${\mathcal{L}}^{{\rm lin}}$ is the loss function $\mathcal{L}$} computed on $\mathcal{D}$} using the linearized model function (2.17) rather than the original model function $f_{\theta_{t}}(\cdot)$. In the linear approximation, the time evolution of the model function becomes linear:
$\frac{{\rm d}}{{\rm d}t}f_{\theta_{t}^{{\rm lin}}}^{{\rm lin}}(x)=-\eta\hat{K}_{\theta_{0}}(x,X_{{\rm train}}^{T})e^{-\eta\hat{K}_{\theta_{0}}t}(F_{\theta_{0},{\rm train}}(0)-Y)$ (2.19)
@@
$f_{\theta_{t}^{{\rm lin}}}^{{\rm lin}}(x)=f_{\theta_{0}}(x)-\hat{K}_{\theta_{0}}(x,X_{{\rm train}}^{T})\hat{K}_{\theta_{0}}^{-1}\left(\mathbbm{1}-e^{-\eta\hat{K}_{\theta_{0}}t}\right)(F_{\theta_{0},{\rm train}}(0)-Y).$ (2.20)
Ref. girardi2024 proves that in the limit of infinite width, if the covariance of $f$ at initialization and the analytic NTK have finite limits $\mathcal{K}_{0}$ and $K$, respectively, the probability distribution of the trained model function $\{f_{\theta_{t}}(x)\}_{x\in{\mathcal{X}}}$ converges in distribution to the Gaussian process with mean and covariance given by (girardi2024, , Corollary 4.16)
@@
Ref. hernandez2024 provides a quantitative version of the convergence that holds at finite width. The results of Ref. hernandez2024 hold uniformly with respect to the training time, and therefore prove that the limits of infinite width and infinite training time commute, and therefore the average of the model function trained for infinite time actually converges to (2.23) in the limit of infinite width.
@@ 3.1 An efficient classical algorithm to estimate the analytic NTK
This is just a matter of computation, and we refer to girardi2024 ; banchi2021measuring . ∎
@@
where we have used the parameter-shift rule to express the derivatives with respect to the parameters as finite differences. From a statistical perspective, the estimator $\tilde{K}$ possesses several desirable properties that make it a reliable tool for approximating the analytic NTK. First, it is an unbiased estimator, meaning that its expected value coincides exactly with the true kernel (see Proposition 1 below). This ensures that, on average, the estimator does not systematically overestimate or underestimate the quantity of interest barankin1949locally ; halmos1946theory . Moreover, because the estimator is constructed as an average over $N$ independent and identically distributed random samples, classical concentration inequalities (such as Hoeffding’s or Bernstein’s inequality) can be applied to provide probabilistic guarantees on the deviation of $\tilde{K}$ from $K$ Tropp_2011 ; talagrand1995missing . We then propose the following classical algorithm to estimate the NTK:
@@ 3.2 Sample complexity
Let $\mathcal{X}$ be the feature space, and for each $x\in\mathcal{X}$ let us consider a parametric quantum circuit as defined in (2.3) with model function as in (2.5). Suppose that Assumption 1 holds true. Let us assume that $\mathbb{E}_{\theta}\,f_{\theta}(x)=0$ for any $x\in\mathcal{X}$. Then for any $x,x^{\prime}\in\mathcal{X}$ and any $\epsilon,\,\delta>0$, the output $\hat{K}(x,x^{\prime})$ of Algorithm 1 with
@@
$\left|\hat{K}(x,x^{\prime})-K(x,x^{\prime})\right|<\epsilon$ (3.4)
@@ 4.2.2 Estimation of the number of elementary operations
where $P_{k}\in{\bf P}_{n}$, $k=1,\ldots,m$. Hence the images of $P_{k}$ under (4.32) can be written as in (LABEL:represent). Furthermore, by the explicit formulas in (LABEL:represent), we have that the number of operations to specify the image is then $O(n^{2})$ (see also (Aaronson_2004, , Section III)). Then for each $\ell=1,\ldots,L$ one gets that by Lemma 2,
@@ 4.3.2 Estimation of the number of elementary operations
Let us now to estimate the number of elementary operations required to compute $\mu_{\infty}(x)$. Recall that by Lemma 3 and Lemma 5, the number of operations required to calculate $f_{\theta}(x)$ and $\nabla_{\theta}f_{\theta}(x)$ is $O(Lmn^{2})$, and $O(L^{2}mn^{2})$ respectively. Furthermore, the complexity of computing $\nabla_{\theta}f_{\theta}(X_{{\rm train}})$ is $O(d_{{\rm train}}L^{2}mn^{2})$. On the other hand, to compute the product $\nabla_{\theta}f_{\theta}(x)\nabla_{\theta}f_{\theta}(X_{{\rm train}}^{T})$ we need to perform $Ld_{{\rm train}}$ products. Hence, to compute $K(x,X_{{\rm train}}^{T})$, we need $O(NL[d_{{\rm train}}Lmn^{2}+d_{{\rm train}}])$ operations. On the other hand, by strassen1969gaussian , the complexity of inverting $K_{{\rm train}}$ is $O(d_{{\rm train}}^{3})$. Hence, the complexity of computing $K_{{\rm train}}^{-1}$ is $O(NL[Lmn^{2}d_{{\rm train}}+d_{{\rm train}}^{2}]+d_{{\rm train}}^{3})$. Therefore, the complexity of computing $\mu_{\infty}(x)$ is then

### v2

@@ Abstract
We propose an efficient classical algorithm to estimate the Neural Tangent Kernel (NTK) associated with a broad class of quantum neural networks. These networks consist of arbitrary unitary operators belonging to the Clifford group interleaved with parametric gates given by the time evolution generated by an arbitrary Hamiltonian belonging to the Pauli group. The proposed algorithm leverages a key insight: the average over the distribution of initialization parameters in the NTK definition can be exactly replaced by an average over just four discrete values, chosen such that the corresponding parametric gates are Clifford operations. This reduction enables an efficient classical simulation of the circuit. Combined with recent results establishing the equivalence between wide quantum neural networks and Gaussian processes [Girardi et al., Comm. Math. Phys. 406, 92 (2025); Melchor Hernandez et al., Ann. Henri Poincaré (2025)], our method enables efficient computation of the expected output of wide, trained quantum neural networks, and therefore shows that such networks cannot achieve quantum advantage.
@@ 1 Introduction
Quantum Machine Learning (QML) is an interdisciplinary field that merges the principles of quantum computing with classical machine learning techniques [9, 25, 21]. One of the core ideas of QML is to harness quantum algorithms and the unique properties of quantum mechanics such as superposition, entanglement, and quantum parallelism to enhance the performance of deep neural models [6]. Quantum neural networks constitute the quantum version of deep neural models. These new models are based on quantum circuits and generate functions given by the expectation values of a quantum observable measured on the output of a quantum circuit made by parametric gates [11, 24]. These parameters are typically optimized by gradient descent, which involves iterative adjustment to minimize a cost function and improve the performance of the quantum circuit in the processing and analysis of data [26]. Several works have focused on the analysis of quantum neural networks, as it is believed that they can combine the computational power of quantum computers with the capabilities of deep learning algorithms [17]. Ref. [16] shows that an exponential quantum speed-up can be obtained via the use of a quantum-enhanced feature space, where each data point is mapped in a non-linear way to a quantum state, and then classified by a linear classifier in a high-dimensional Hilbert space [13]. Nevertheless, a significant disadvantage lies in the need to determine the appropriate parameters to configure the quantum circuit beforehand, and it is not yet clear whether these parameters can be effectively obtained using a variational technique [8].
@@
A rigorous mathematical characterization of the training dynamics of quantum neural networks becomes possible in the limit of infinite width. Ref. [2] proves that in the limit of infinite width with the depth kept fixed, trained quantum neural networks with constant depth operate in the lazy regime (i.e., the maximum amount by which the training can change a parameter tends to zero) and are capable of perfectly fitting the training examples. However, such a regime is not interesting for quantum advantage, since at finite depth, the past light-cone of each measured qubit has a finite size and the expectation value of the measured observable can be efficiently estimated on a classical computer. The recent works [11, 20] consider the joint limit of infinite width and depth. Under the hypothesis that the depth grows at most logarithmically with respect to the number of qubits, they prove that the probability distribution of the trained model function converges in distribution to a Gaussian process. The key element of the proof is showing that the training happens in the lazy regime, and therefore the dependence of the model function on the parameters can be approximated by its linearized version near the initialization values. Consequently, in the limit the model becomes linear and the training has an analytic solution whose probability distribution is Gaussian with analytically computable mean and covariance. Such a linear model is governed by a kernel called the Neural Tangent Kernel (NTK), which has been investigated in several recent studies [25, 27, 15, 32, 23].
@@
Let us recall that, given a general parametric model $f_{\theta}:\mathcal{X}\rightarrow\mathbb{R}$, with $\theta\in\mathbb{R}^{L}$, where $\mathcal{X}$ is the feature space, the empirical NTK is defined as [14]:
@@ 1.1 Our contribution
The key idea of the algorithm is that, since $f_{\theta}(x)$ is a quadratic polynomial in the entries of the unitary matrix associated to each gate and the analytic NTK is quadratic in $f$, the expectation value with respect to the uniform distribution of the parameters can be replaced by the expectation value with respect to the uniform distribution on the set $\left\{0,\frac{\pi}{2},\pi,\frac{3\pi}{2}\right\}$. For such values of the parameters, the parametric gates belong to the Clifford group, and since also the nonparametric gates belong to the Clifford group by hypothesis, the whole $U_{x,\theta}$ belongs to the Clifford group. Therefore, the analytic NTK can be computed via the efficient algorithm for Clifford quantum circuits [1].
Let us now state informally our second contribution. The average of the Gaussian process associated to a trained quantum neural network has a limit $\mu_{\infty}:\mathcal{X}\to\mathbb{R}$ for infinite training time. Such a limit depends only on the NTK of the network and on the training data (see subsection 2.4 and [11, 20] for the details). We can now leverage the algorithm of Theorem 1 to estimate $\mu_{\infty}$:
@@
    The output $\tilde{\mu}_{\infty}$ of the algorithm above satisfies
@@
Remark 2.
Theorem 2 states that sample and time complexities related to the classical algorithm for estimating $\mu_{\infty}(x)$ are polynomial. However, we do not believe that our exponents are optimal. Our result needs a time complexity scaling as $d_{\rm train}^{6}$ (see (1.6) and (1.9)). For realistic values of $d_{\rm train}$ (e.g., the MNIST dataset of handwritten digits includes about $10^{5}$ examples), such a scaling would be prohibitive. We are confident that our upper bound to the time complexity is pessimistic and that numerical experiments, which we plan to address in a future work, may reveal a better scaling.
Our work conceptually aligns with a recent series of results [3, 22, 33, 18, 7] demonstrating that a wide class of variational quantum circuits can be simulated efficiently by classical computers. In particular, Ref. [3] shows that, if each layer of the circuit contains random iid one-qubit gates sampled from the Haar distribution applied to all the qubits, then, regardless of how the two-qubit gates are placed, the untrained circuit can be simulated efficiently with high probability. Our results imply that wide quantum neural networks with logarithmic depth trained on supervised-learning problems with the input encoded with Clifford gates can be simulated efficiently by classical computers.
@@ 2.2 The Pauli group and the Clifford group
For details about properties of these groups, see for instance [19].
@@ 2.4 The neural tangent kernel
In what follows, we assume the following:
Assumption 1.
The parameter vector $\theta$ is initialized by independently sampling each entry from the uniform distribution on $0,2\pi)$.
Here, we note that the parameters are optimized via gradient descent with respect to the quadratic cost function $\mathcal{L}$} as follows:
where $\mathcal{L}$} is given by
and $\eta>0$ is the learning rate. Given any function $g:\mathcal{X}\to\mathbb{R}$, we will often use the following notation:
@@
From the gradient-flow equation (2.8) and the chain rule, one gets that
@@
where $\nabla_{\theta}f_{\theta_{t}}(X_{{\rm train}}^{T})$ denotes the gradient of $f_{\theta_{t}}(X_{{\rm train}}^{T})$ with respect to $\theta$. Recall that T is the transposition operator. Let us notice that (2.15) can be written as
$\displaystyle\begin{cases}\frac{{\rm d}\theta_{t}}{{\rm d}t}=-\eta\nabla_{\theta}f_{\theta_{t}}(X_{{\rm train}}^{T})(F_{\theta_{t},{\rm train}}-Y),\\ \frac{{\rm d}}{{\rm d}t}f_{\theta_{t}}(x)=-\eta\hat{K}_{\theta_{t}}(x,X_{{\rm train}}^{T})(F_{\theta_{t},{\rm train}}-Y).\end{cases}$ (2.16)
If we assume that the NTK does not change during training, i.e., that $\hat{K}_{\theta_{t}}=\hat{K}_{\theta_{0}}$, then the time evolution of the model function becomes linear:
$\frac{{\rm d}}{{\rm d}t}f_{\theta_{t}^{{\rm lin}}}^{{\rm lin}}(x)=-\eta\hat{K}_{\theta_{0}}(x,X_{{\rm train}}^{T})e^{-\eta\hat{K}_{\theta_{0}}t}(F_{\theta_{0},{\rm train}}-Y)$ (2.17)
@@
$f_{\theta_{t}^{{\rm lin}}}^{{\rm lin}}(x)=f_{\theta_{0}}(x)-\hat{K}_{\theta_{0}}(x,X_{{\rm train}}^{T})\hat{K}_{\theta_{0}}^{-1}\left(\mathbbm{1}-e^{-\eta\hat{K}_{\theta_{0}}t}\right)(F_{\theta_{0},{\rm train}}-Y).$ (2.18)
Ref. [11] proves that this assumption holds: in the limit of infinite width with the depth growing at most logarithmically with the width, if covariance of $f$ at initialization and the analytic NTK have finite limits $\mathcal{K}_{0}$ and $K$, respectively, then the empirical NTK stays close to its value at initialization throughout the whole training. Moreover, the value of the NTK at initialization is with high probability close to its average with respect to the random initialization of the parameters [11, Theorem 4.7]. As a consequence, the probability distribution of the trained model function $\{f_{\theta_{t}}(x)\}_{x\in{\mathcal{X}}}$ converges in distribution to the Gaussian process with mean and covariance given by [11, Corollary 4.9]:
@@
Ref. [20] provides a quantitative version of the convergence that holds at finite width. The results of Ref. [20] hold uniformly with respect to the training time, and therefore prove that the limits of infinite width and infinite training time commute, and therefore the average of the model function trained for infinite time actually converges to (2.21) in the limit of infinite width.
@@ 3.1 An efficient classical algorithm to estimate the analytic NTK
This is just a matter of computation, and we refer to [11, 4]. ∎
@@
where we have used the parameter-shift rule to express the derivatives with respect to the parameters as finite differences. From a statistical perspective, the estimator $\tilde{K}$ possesses several desirable properties that make it a reliable tool for approximating the analytic NTK. First, it is an unbiased estimator, meaning that its expected value coincides exactly with the true kernel (see Proposition 1 below). This ensures that, on average, the estimator does not systematically overestimate or underestimate the quantity of interest [5, 12]. Moreover, because the estimator is constructed as an average over $N$ independent and identically distributed random samples, classical concentration inequalities (such as Hoeffding’s or Bernstein’s inequality) can be applied to provide probabilistic guarantees on the deviation of $\tilde{K}$ from $K$ [30, 29]. We then propose the following classical algorithm to estimate the NTK:
@@ 3.2 Sample complexity
Let $\mathcal{X}$ be the feature space, and for each $x\in\mathcal{X}$ let us consider a parametric quantum circuit as defined in (2.3) with model function as in (2.5). Suppose that Assumption 1 holds true. Let us assume that $\mathbb{E}_{\theta}\,f_{\theta}(x)=0$ for any $x\in\mathcal{X}$. Then for any $x,x^{\prime}\in\mathcal{X}$ and any $\epsilon,\,\delta>0$, the output $\tilde{K}(x,x^{\prime})$ of Algorithm 1 with
@@
$\left|\tilde{K}(x,x^{\prime})-K(x,x^{\prime})\right|<\epsilon$ (3.4)
@@
Remark 4.
We stress that the algorithm itself does not require any assumptions on the width or depth of the network. Such assumptions are only needed to guarantee that $\mu_{\infty}$ is close to the mean of the trained model function.
@@ 4.2.2 Estimation of the number of elementary operations
where $P_{k}\in{\bf P}_{n}$, $k=1,\ldots,m$. Hence the images of $P_{k}$ under (4.32) can be written as in (LABEL:represent). Furthermore, by the explicit formulas in (LABEL:represent), we have that the number of operations to specify the image is then $O(n^{2})$ (see also [1, Section III]). Then for each $\ell=1,\ldots,L$ one gets that by Lemma 2,
@@ 4.3.2 Estimation of the number of elementary operations
Let us now to estimate the number of elementary operations required to compute $\mu_{\infty}(x)$. Recall that by Lemma 3 and Lemma 5, the number of operations required to calculate $f_{\theta}(x)$ and $\nabla_{\theta}f_{\theta}(x)$ is $O(Lmn^{2})$, and $O(L^{2}mn^{2})$ respectively. Furthermore, the complexity of computing $\nabla_{\theta}f_{\theta}(X_{{\rm train}})$ is $O(d_{{\rm train}}L^{2}mn^{2})$. On the other hand, to compute the product $\nabla_{\theta}f_{\theta}(x)\nabla_{\theta}f_{\theta}(X_{{\rm train}}^{T})$ we need to perform $Ld_{{\rm train}}$ products. Hence, to compute $K(x,X_{{\rm train}}^{T})$, we need $O(NL[d_{{\rm train}}Lmn^{2}+d_{{\rm train}}])$ operations. On the other hand, by [28], the complexity of inverting $K_{{\rm train}}$ is $O(d_{{\rm train}}^{3})$. Hence, the complexity of computing $K_{{\rm train}}^{-1}$ is $O(NL[Lmn^{2}d_{{\rm train}}+d_{{\rm train}}^{2}]+d_{{\rm train}}^{3})$. Therefore, the complexity of computing $\mu_{\infty}(x)$ is then